OpenStudy (anonymous):

If A and B are subsets of a set X, prove that A\B = A intersects complement of B.

OpenStudy (anonymous):

Ok this is my guess ... feel free to critique Suppose \(A\subseteq X\) and \(B \subseteq X\) such that every element of A and B are elements of X Suppose \(y \in A\) and \(y \notin B\) so\( y \in A−B\) Hence \(y \in A\) and \(y \in B^c\) so \(y \in A \cap B^c\) Therefore \(A−B=A \cap B^c\)

OpenStudy (anonymous):

Im just wondering if Im missing a line in there Like maybe im somehow supposed to say that B^c is also a subset of X but im not sure that every element of B^c is an element of X

OpenStudy (freckles):

usually to prove an equality you prove one is a subset of the other by supposing there is an element in A-B and then showing it is also in A and not B. Then you suppose there is an element in A and not B and show it is also in A-B.

OpenStudy (anonymous):

ohhhh yeaaaaaaa

OpenStudy (freckles):

Like I would rewrite what you have so far as this: We want to prove \[A-B \underline{\subset }A \cap B^c \\ \text{ Let } y=A-B \\ \text{ then by definition of } A-B \\ y \in A \text{ and } y \not \in B \\ \text{ so that implies } \\ y \in A \text{ and } y \in B^c \\ \text{ therefore } y \in A \cap B^c\] now you show the other direction show \[A \cap B^c \underline{\subset}A-B\] first line: \[\text{ Suppose } x \in A \cap B^c \] -- though I guess you could say some of the things we are saying seem redundant :p

OpenStudy (freckles):

you can also prove this using a truth table

OpenStudy (anonymous):

ya proving sometimes tend to be redundant but i prefer more steps then less cuz its easier to follow Otherwise im like .... waitttt how did they get there??? Anyways thanks :)

OpenStudy (freckles):

and you still have to go that other direction

OpenStudy (anonymous):

Ya the other direction is pretty much similar

OpenStudy (freckles):

oh wait you know what? I was going to prove this by truth table and I just remember that I have always defined A-B as A and not B :p so truth table with not work since I used that as definition

OpenStudy (freckles):

it is actually kinda weird to prove it non truth table wise because we are still saying A-B means A and not B

OpenStudy (anonymous):

ik but i remember when i used to prove this crap we wld basically just do that ANDDD A-B mean the element is in A and Not In B

OpenStudy (freckles):

oops I didn't realize this wasn't your question

OpenStudy (anonymous):

lol yea just reviewing this stuff cuz im bored -.- Or rather this a form of procrastination I prefer studying leisurely than studying stuff I must know Got a huge exam and this is a way of just forgetting abt it ;) Kinda sad lol

OpenStudy (zzr0ck3r):

just change your statements to iff and you are done ;) \(x\in A-B \iff x\in A \text{ and } x\notin B \iff x\in A \text{ and } x\in B^C\iff x\in A\cap B^c\)